OpenStudy (anonymous):

MEDAL AND FAN! Simplify Square Root of -48 . A) -4 square root of 3 B) 4 square root of negative 3 C) 4i square root of 3 D) 4 square root of 3 i

OpenStudy (anonymous):

@mathmate could you please help me?

OpenStudy (anonymous):

I don't think it is A or B since we are learning about i

OpenStudy (gabebae):

I know this

OpenStudy (gabebae):

It's A You can't have a negative

OpenStudy (gabebae):

no negatives under a square root

OpenStudy (anonymous):

c and d do not have negatives

OpenStudy (mathmate):

Hints: 1. The square-root of a negative numerical quantity has an "i" attached to it, as you suggest, but for a different reason. By definition, y=sqrt(x) means y^2=x. Since i^2=-1, so y^2 becomes negative if y contains i as a factor. Example: \((4i)^2 = 4^2 i^2 =16(-1)=-16\) (note brackets around 4i. 2. when simplifying sqrt(x) when x is composite (i.e. consists of numerous factors), the work can be simplified by first factoring x. For example, \(\sqrt{96}=\sqrt{2^5*3}=\sqrt{2^4*2*3}=\sqrt{2^4}*\sqrt{2*3}=2^2\sqrt{6}=4\sqrt 6\) With the two hints understood, you will have no problem solving the posted problem.

OpenStudy (anonymous):

So I am supposed to square and multiply then simplify okie ^_^

OpenStudy (gabebae):

oh i'm sorry I didn't pay attention to the answers

OpenStudy (mathmate):

If you combine the two hints, you would get \(\sqrt{-96}=4\sqrt 6 i\) Check: \((4\sqrt 6 i)^2 =4^2(\sqrt 6)^2i^2=16(6)(-1)=-96, so that's correct!\)